OpenStudy (anonymous):

write in standard form an equation of each line slope=-3 through (4,0)

OpenStudy (fanduekisses):

Use the point slope formula

OpenStudy (fanduekisses):

and then get it in this form: Ax + Bx = C

OpenStudy (fanduekisses):

The point slope form: \[y-y_{1}=m(x-x_{1})\] so plug in what you have, m= -3 (slope) y1=0 and x1=4

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 4}}\quad ,&{\color{blue}{ 0}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= -3 \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and move everything to the left}\\ \qquad \uparrow\\ \textit{point-slope form}\)